OpenStudy (anonymous):

36π cm³ of air is used to fill up a special balloon. Find the surface area (in terms of π) and the diameter of the balloon.

OpenStudy (anonymous):

We can assume that the special balloon is sphere-like and use the formula: Volume of a sphere 

OpenStudy (anonymous):

Volume of a sphere=\[\frac{ 4 }{ 3 } \pi r^3\] the volume is 36π so you just put it in the formula: \[36\pi=\frac{ 4 }{ 3 }\pi r^3\] can you solve this or do you need help?